OpenStudy (anonymous):

Quick for a medal! If f(x)=x-5 AND g(x)=x^2+3, find f[g(-2)]

OpenStudy (anonymous):

\(g(-2)=(-2)^2+3=4+3=7\) all that is left is for you to find \(f(7)=7-5\)

OpenStudy (anonymous):

so it's 2?